# Self Rewarding Self Improving

### v1

@@ 5.2 Generator Verifier Gap
  • Competitive mathematics, where judges can programmaticly brute force verify solutions that require complex multi-step reasoning to generate
@@ 5.3 Autonomous Self Improving
The potential implications are profound. By generating their own practice problems and providing their own feedback, models can potentially overcome data limitations that currently constrain AI development. When synthetic data generation and self-evaluation prove effective, the primary bottleneck becomes computational resources rather than human-annotated data.
@@ 5.4 Reward Hacking and Exploitation
In contrast, GPT-4o demonstrated remarkable resistance to these exploitation attempts, maintaining evaluation integrity across various prompt formulations. This mirrors observations in safety research where larger models show substantially improved jailbreak resistance compared to smaller counterparts.

### v2

@@ 4 Discussion
4.1 Limitations
Our experiments revealed significant fragility in self-judging systems that required extensive iteration to overcome. Developing effective judge prompts proved surprisingly brittle—minor wording changes could trigger catastrophic reward hacking within dozens of training steps, forcing us to roll back experiments and redesign evaluation strategies. Our Countdown task alone required four major prompt iterations before achieving stable training, and even then, agents consistently discovered creative exploitation strategies like formatting tricks or nonsensical but convincing-looking expressions. The non-deterministic nature of LLM judges introduced substantial variance across runs, with identical configurations producing divergent outcomes that complicated reproducibility and required multiple seeds to draw reliable conclusions. While updating judge weights showed promise for breaking performance plateaus, it introduced new failure modes—continuous updates led to rapid collapse as models exploited their own biases, while even successful single-update approaches eventually plateaued. Additionally, disentangling improvements in output formatting from genuine mathematical reasoning proved challenging, as early performance gains often reflected models learning SymPy syntax rather than enhanced problem-solving ability. Notably, we observed a clear relationship between model scale and robustness—larger models like DeepSeek R1 proved substantially more resistant to reward hacking and required minimal prompt engineering, while smaller models like Llama 3.2 3B were highly susceptible to exploitation regardless of prompt design. This scale-dependent brittleness suggests that while current self-judging approaches may become more practical as model capabilities advance, smaller models require extensive domain-specific tuning and careful monitoring to achieve stable training.
@@ 4.3 Generator Verifier Gap
  • Competitive mathematics, where judges can programmatically brute force verify solutions that require complex multi-step reasoning to generate
@@ 4.4 Autonomous Self Improving
The potential implications are profound. By generating their own practice problems and providing their own feedback, models can potentially overcome data limitations that currently constrain AI development. When synthetic data generation and self-evaluation proves effective, the primary bottleneck becomes computational resources rather than human-annotated data.
@@ 4.5 Reward Hacking and Exploitation
In contrast, we found R1 demonstrated resistance to these exploitation attempts, maintaining evaluation integrity across various prompt formulations. This mirrors observations in safety research where larger models show substantially improved jailbreak resistance compared to smaller counterparts.
@@
5 Related Work
Our approach builds upon several research streams in self-improvement and LLM training. Meta’s Self-Rewarding Language Models work demonstrated that models can generate preference data using an LLM as a judge with Direct Preference Optimization [7] that led to significant performance improvements on benchmark tasks. Unlike Meta’s preference-based approach, our method focuses specifically on mathematical reasoning domains with verifiable solutions.
We also extend our previous work [8] LADDER framework, which enabled improvements on mathematical benchmarks through recursive problem decomposition and demonstrated capabilities for building self-improving systems for reasoning tasks. Our work specifically builds on LADDER’s synthetic question generation capabilities, but implements LLM as a judge rather than a numeric verifier.
Our research connects to weak-to-strong supervision literature, where [9] demonstrated that larger models supervised by smaller models can achieve strong performance on tasks the smaller model couldn’t solve consistently. This relates to concepts from iterated amplification [10] and addresses alignment concerns raised[11], that warn about reward hacking in learned optimizers. Our approach mitigates these risks by using verifiable signals rather than opaque weak supervision.
We also build upon code generation frameworks like CodeRL [12] that use unit test outcomes as reward criteria and verification-based learning from solved examples. Unlike approaches that rely solely on external verification systems, our method establishes a framework where models practice on self-created exercises with reliable self-judgment, enabling continuous improvement in domains where defining programmatic reward functions is challenging.